OpenStudy (anonymous):

Ten students begin college at the same time. The probability of graduating in four years is 63%. Which expanded expression shows the first and last terms of the expression used to find the probability that at least six will graduate in four years?

OpenStudy (anonymous):

(106)×(0.63)6×(0.37)4+...+(1010)×(0.63)10×(0.37)∘=0.706 (105)×(0.63)5×(0.37)5+...+(1010)×(0.63)10×(0.37)∘=0.88 (100)×(0.63)∘×(0.37)10+...+(106)×(0.63)6×(0.37)4=0.540 (100)×(0.63)∘×(0.37)10+...+(105)×(0.63)5×(0.37)5=0.293

OpenStudy (anonymous):

@Black_ninja123